OpenStudy (tishockjo):

A six-sided number cube with faces numbered 1, 2, 3, 4, 5, and 6 is rolled. What is the probability of rolling a 6 or a number less than 3? A.1/8 B.1/6 C.1/3 D.1/2

OpenStudy (anonymous):

{1,2,6} FROM {1,2,3,4,5,6} THUS REQD PROB = 3/6 =1/2

OpenStudy (aric200):

Why did he ask it twice

OpenStudy (tishockjo):

@aric200 sr...I did it by mistake

OpenStudy (tishockjo):

:)

OpenStudy (anonymous):

P(6 or <3) = 1/6 + 2/6 =3/6 =1/2 hope that helps.

OpenStudy (aric200):

A little late @JoeMGachau